# The GenAI Generation: Student Views of Awareness, Preparedness, and Concern

### v1

@@ 5 Conclusion
The GenAI Generation, defined as students in their formative educational years during the rise of GenAI, face a rapidly evolving landscape shaped by both opportunity and uncertainty. While many are excited about GenAI’s potential, they also express concerns around preparedness, ethics, and long-term implications. Findings show that students with greater exposure to GenAI, particularly graduate computer science students, report higher confidence and optimism about their future careers. However, such structured exposure is not consistent across disciplines, revealing an institutional gap that must be addressed to ensure equitable support for all students navigating this technological shift.
@@
The GenAI Generation is distinct in its developmental context as students are not only adapting to new technologies but doing so without established norms or clear guidelines. They are the early adopters of tools like ChatGPT, learning to balance traditional skills with new competencies required to interact with GenAI effectively. Their adaptability is frequently tested while grappling with ethical questions, workforce changes, and broader societal implications.
Students want more than just exposure: they seek guidance on using GenAI tools responsibly, critical thinking around AI-generated outputs, and a robust understanding of ethical considerations. The GenAI Generation is asking for a holistic approach that not only enhances technical skills but also nurtures critical engagement and ethical reasoning. It is critical to closely monitor and derive clarity on the potential impacts of these tools, developing a deeper understanding of cognition and ethical implications.
@@
The GenAI Generation is at a pivotal moment. As the first cohort to learn and grow in a GenAI-centric world, their experiences will set the tone for future generations. The data highlights the need for GenAI literacy efforts that span disciplines and prepare both students and faculty to engage with these tools in meaningful ways. The voices of these students are clear: they want to be part of a future where GenAI is a tool for growth, not a source of uncertainty, and where their education equips them to use GenAI as a force for positive change.
Acknowledgements
We would like to thank the students who took the time to respond to our survey, Dr. Joseph Ludlum and Dr. Loraine Phillips from the Office of Academic Effectiveness at Georgia Tech for their support in distributing the Fall 2023 cohort survey, and Dr. Thomas Ploetz for the continued encouragement throughout the process.

### v2

@@ 5 Conclusion
This study contributes to the growing discourse on generative AI in education by examining how students are navigating both opportunity and uncertainty without consistent institutional guidance. It introduces the term GenAI Generation to describe those in their formative educational years during the rise of GenAI, students whose academic development is unfolding alongside the rapid adoption of these tools. Through analysis of over 250 survey responses, the findings reveal not only patterns of perceived readiness and ethical concern, but also a broader psychosocial landscape and pedagogical gap. Students are engaging with GenAI amid unclear norms, forming academic identities and career expectations as institutions lag behind. These insights offer timely direction for educators and policymakers seeking to design inclusive, developmentally aligned, and ethically grounded approaches to GenAI integration.
The GenAI Generation faces a rapidly evolving landscape shaped by both excitement about GenAI’s potential and concern about preparedness, ethics, and long-term implications. This study reveals a growing divide in student confidence and readiness, shaped more by exposure and self-directed integration of GenAI than by formal curriculum. However, structured incorporation is inconsistent across disciplines, highlighting a broader institutional gap that must be addressed to ensure equitable support for all students navigating this technological shift. The divide is no longer simply between STEM and non-STEM fields, but between those with and without access to meaningful engagement with emerging technologies.
@@
The GenAI Generation is distinct in its developmental context. These students are not only adapting to powerful new tools, but doing so without established norms or guidance. As early adopters, they must balance traditional skills with new competencies while grappling with ethical dilemmas, shifting workforce expectations, and unknown cognitive impacts. Their adaptability is being tested in real time.
Students want more than access, they are asking for structure. They seek guidance on using GenAI responsibly, developing critical thinking around AI-generated outputs, and building a strong foundation in ethical reasoning. Their feedback calls for a holistic approach that prepares them not only to use GenAI, but to question and shape it.
@@
As the first cohort to learn and grow in a GenAI-centric world, the experiences of the GenAI Generation will set the tone for future educational trajectories. Their voices point to an urgent need for GenAI literacy that spans disciplines and prepares both students and faculty to engage with these tools thoughtfully and equitably. They envision a future where GenAI is not a source of confusion or inequity, but a tool for growth—integrated into education in ways that are ethical, empowering, and human-centered.
6 List of abbreviations
Abbreviation Full Form
GenAI Generative Artificial Intelligence
AI Artificial Intelligence
IRB Institutional Review Board
OMSCS Online Master of Science in Computer Science
MSCS Master of Science in Computer Science
OAE [Institution Name Redacted for Review] Office of Academic Effectiveness
STEM Science, Technology, Engineering, and Mathematics
CS Computer Science (used in tables/figures for grouping)
7 Future Considerations
To better understand the evolving impact of GenAI on student development and educational outcomes, future research should consider longitudinal studies that track student perceptions, readiness, and concerns over time. As GenAI tools rapidly advance, students’ experiences and attitudes will likely shift in response to both technological innovation and institutional adaptation. Capturing these changes across semesters or academic years would provide valuable insight into how exposure, policy, and pedagogy shape preparedness and perception in the long term.
In addition to tracking perception, researchers should examine the cognitive consequences of increased reliance on GenAI tools, such as ChatGPT, for learning and problem-solving. Cognitive offloading may free students to focus on higher-order thinking, but it also raises questions about memory retention, deep learning, and critical reasoning. Understanding how sustained use of GenAI affects cognitive engagement, information retention, and academic identity will be critical to developing pedagogical strategies that support—not replace—core learning processes.
There is also a need for targeted dissemination and outreach to disciplines beyond computer science. This study found that students in non-computer science fields, particularly those in the ”Other” category, demonstrated strong emotional responses and a desire for inclusion in GenAI discourse, despite limited curricular exposure. Expanding GenAI literacy and integration into these areas can help bridge the readiness gap and empower students across all domains to engage critically and confidently with this transformative technology.
8 Declarations
8.1 Funding
This study was conducted without external funding. No funding agency was involved in the design, data collection, analysis, or publication of this research.
8.2 Funding Statement
8.3 Consent to Publish declaration
Not applicable
8.4 Consent to Participate
This research study was Institutional Review Board (IRB) approved. Respondents provided informed consent.
8.5 Ethics declaration
This research study was Institutional Review Board (IRB) approved. The study collected students’ perceptions of their future career preparedness in light of the potential impact of generative artificial intelligence (GenAI) through a survey. The survey only collected responses for students located within the United States, and 18 years of age or older as per IRB instruction. Respondents provided informed consent. Survey questions were designed to protect participant identity. Participation was completely voluntary, and had the ability to opt out at any time. No benefit or compensation was provided. We complied with any applicable laws and regulations regarding confidentiality. The IRB and Office of Human Research Protections reserved the right to look at study records at any time. Contacts to the author and Office of Research Integrity Assurance were provided.